# Physics-Aware Semi-Supervised Underwater Image Enhancement

### v2

@@ 5.2 Evaluation on Transmission Estimation and Color Restoration
The quantitative evaluation of our methods and eight baselines for transmission estimation and color restoration is reported in Table 5.2. It can be seen that our method achieved the superior, or at least the comparable, performance to that of the eight baselines on four different subsets of Test-R53. Specifically, our method outperformed the baselines with a large margin on both the Michmoret and Katzaa subsets no matter which metric was considered, while it performed comparably to the best baseline on the Nachsholim and Satil subsets. It should be noted that none of these methods produced a PCC value higher than 0.12 on the Satil subset. The inferior results may be attributed to the special scenes contained in this subset [25]. As shown in Fig. 5, our method is able to perform transmission estimation and color restoration by directly learning from the limited real-world underwater images.
@@
Table 2: The quantitative evaluation of eight baselines and our method on the four subsets of Test-R53, where PCC and $\bar{\psi}$ are used for transmission estimation and color restoration respectively.
@@
DHL -0.07 32.10 0.10 35.50 0.39 35.91 0.11 34.82
HE - 35.44 - 35.43 - 34.37 - 36.33
Retinex - 31.44 - 34.37 - 33.29 - 35.31
WaterNet - 25.40 - 25.70 - 25.80 - 28.31
UColor - 15.55 - 13.94 - 17.21 - 21.43
Ours 0.71 11.51 0.48 8.75 0.18 19.59 0.04 13.46
Table 3: The ablation study on Test-U80 and the Michmorest subset in Test-R53.
Method Test-U80 Michmoret
@@
Simplified 22.16 89.57 -0.15
w/ RCT 22.97 89.93 0.58
w/ Bi-Sup 22.51 88.94 0.68
w/ Semi-Sup 22.82 88.87 0.71
5.3 Ablation Study
To examine the effect of each component of our method, including the RCT, bi-directional supervised scheme and semi-supervised learning framework, we conducted an ablation study. For simplicity, we only utilized Test-U80 and the Michmoret subset in Test-R53. The PA-UIENet without the RCT together with the forward-enhancement supervised learning module was crafted as a simplified version. On top this version, we derived three additional variants by inserting the RCT (w/ RCT) and then replacing the forward-enhancement supervised module by the bi-directional supervised scheme (w/ Bi-Sup) or the semi-supervised learning framework (w/ Semi-Sup). As shown in Table 5.2, the simplified version outperformed all the baselines in terms of the PSNR and SSIM metrics. However, this variant did not estimate transmission maps well, which may be attributed to the overfitting problem resulted from limited training data. When the RCT was used, the performance of image enhancement and transmission estimation was boosted. It was indicated that the adaptability of the red channel increased the robustness and generalization of the model. Though the utilization of the bi-directional-supervised learning scheme or the semi-supervised learning framework had a negative effect on the PSNR and SSIM values, the effectiveness of transmission estimation continuously increased. It is suggested that the use of the backward-degradation learning module or the unsupervised learning scheme alleviates the overfitting problem.
6 Conclusion
@@
In this paper, we developed a novel Physics-Aware Dual-Stream Underwater Image Enhancement Network, namely, PA-UIENet, which explicitly estimates the degradation parameters of the IFM, for the UIE task. To overcome the challenge of insufficient data, we also adopted an IFM-inspired semi-supervised learning framework, comprising a bi-directional supervised scheme and an unsupervised scheme. Due to the complementary action of the two schemes, the PA-UIENet can be better trained using this framework with both the labeled and unlabeled real-world images while the generalization of the model is stronger, compared to supervised and unsupervised methods. To our knowledge, this study made the first attempt to jointly exploit the physics-aware deep network and the IFM-inspired semi-supervised learning technique for the UIE task. The proposed method performed better than, or at least comparably to, eight baselines on five real-world underwater testing sets in different evaluations. The promising results should be due to the fact that our method is able to not only estimate degradation parameters but also learn the characteristics of diverse underwater scenes.

### v3

@@ V-B Transmission Estimation and Color Restoration
The quantitative evaluation of our methods and four prior-based baselines for transmission estimation and color restoration is reported in Table III. It can be seen that our method achieved the superior, or at least the comparable, performance to that of the four prior-based baselines on four different subsets of Test-R53. Specifically, our method outperformed the baselines with a large margin on both the Michmoret and Katzaa subsets, no matter which metric was considered. It should be noted that none of these methods produced a PCC value higher than 0.12 on the Satil subset. The inferior results may be attributed to the special scenes contained in this subset [35]. As shown in Fig. 10, our method was able to perform transmission estimation and color restoration by directly learning from the limited real-world underwater images.
@@
TABLE III: The quantitative evaluation of four prior-based baselines and our method on the four subsets of Test-R53, where PCC and $\bar{\psi}$ are used for transmission estimation and color restoration, respectively.
@@
DHL -0.07 32.10 0.10 35.50 0.36 35.92 0.11 34.82
Ours 0.61 11.68 0.32 9.47 -0.38 16.87 0.02 10.36
V-C Ablation Study
To investigate the effect of different components of the PATS-UIENet, we conducted a series of ablation experiments. For simplicity, we only utilized Test-U80 and the Michmorest subset in Test-R53.
V-C1 Effect of the Semi-supervised Learning Framework
To further validate the effect of semi-supervised learning framework, we provide a quantitative comparison among the supervised learning, bi-directional supervised learning, and semi-supervised learning frameworks. As reported in Table IV, either the supervised learning framework or the bi-directional supervised learning framework normally generated the lower metric values than those produced by the proposed semi-supervised learning framework, in terms of the PSNR, SSIM, PCC and $\bar{\psi}$ metrics. It has been suggested that our network could be better trained using the semi-supervised learning framework with both the labeled and unlabeled real-world images, which improved the enhancement performance.
TABLE IV: The ablation study of different learning frameworks using Test-U80 and the Michmorest subset in Test-R53.
Learning Framework Test-U80 Michmoret
@@ V-C1 Effect of the Semi-supervised Learning Framework
Sup 21.09 86.72 0.51 17.86
Bi-sup 22.06 88.95 0.65 15.59
Semi-sup (Ours) 23.01 89.89 0.61 11.68
V-C2 Effect of the RCT and the RCM
To examine the effect of different modules on the performance of our PATS-UIENet, we removed the RCT, the RCM or both of them. In this case, we derived three variants of the PATS-UIENet: one with both RCT and RCM removed, one with only RCT removed and one with only RCM removed, denoted as “Simplified”, “w/o RCT” and “w/o RCM”, respectively. The results produced by these variants and the PATS-UIENet are reported in Table V. It can be observed that the PATS-UIENet with the entire components normally achieved the best performance in terms of the PSNR, SSIM, PCC and $\bar{\psi}$ metrics. Although the utilization of the RCT had a negative effect on the PSNR value (observed by comparing “Simplified” with “w/o RCM”), the performance of transmission estimation was increased. It was indicated that the adaptability of the red channel increased the robustness and generalization of the model trained.
TABLE V: The ablation study of different modules on Test-U80 and the Michmorest subset in Test-R53.
Variant Test-U80 Michmoret
[…]
Simplified 22.18 88.56 0.59 13.35
w/o RCT 22.71 89.43 0.60 11.42
w/o RCM 21.84 88.98 0.61 14.05
PATS-UIENet 23.01 89.89 0.61 11.68
V-C3 Effect of Different Network Depth
To evaluate the impact of the network depth on the performance of our PATS-UIENet, we changed the number of layers of the network from 4 to 3 and 5, individually. In total, we derived two variants of the PATS-UIENet with regard to different network depths. In terms of both the variants, the numbers of filters at each block were set to 64, 128, 256, 256, 128, 64 and 64 and 64, 128, 256, 512, 1024, 1024, 512, 256, 128, 64 and 64, respectively. As shown in Table VI, the proposed PATS-UIENet which comprised four layers produced the more promising results compared with the two variants.
TABLE VI: The ablation study of different network depths on Test-U80 and the Michmorest subset in Test-R53.
Network Depth Test-U80 Michmoret
[…]
3 Layers 22.43 89.40 0.61 13.29
5 Layers 22.57 89.45 0.62 9.41
4 Layers (Ours) 23.01 89.89 0.61 11.68
V-C4 Effect of the Revised IFM
To further investigate the effect of the revised IFM [2] on UIE, we used the outputs of the B-Stream and A-Stream of the proposed PATS-UIENet as the transmission map and the ambient light data, respectively, required by the original IFM [1], as shown in Eq. (1). In this case, the results obtained using the original IFM are compared with those produced by the revised IFM in Table VII. As can be seen, the revised IFM achieved the superior performance to the original IFM in terms of the PSNR, SSIM and $\bar{\psi}$ metrics. It is indicated that the revised IFM models the underwater degradation process better than the orignal IFM.
TABLE VII: The comparison between the original IFM [1] and the revised IFM [2] on Test-U80 and the Michmorest subset in Test-R53.
Physical Model Test-U80 Michmoret
[…]
Original IFM [1] 21.93 88.24 0.69 16.24
Revised IFM [2] 23.01 89.89 0.61 11.68
VI Conclusion
@@
In this paper, we introduced a novel Physics-Aware Triple-Stream Underwater Image Enhancement Network, namely, PATS-UIENet, which explicitly estimates the degradation parameters of the revised IFM, for the UIE task. To overcome the challenge of insufficient data, we also adopted an IFM-inspired semi-supervised learning framework, comprising a bi-directional supervised scheme and an unsupervised scheme. Due to the complementary action of the two schemes, the PATS-UIENet can be better trained using this framework with both the labeled and unlabeled real-world images while the generalization of the model is stronger, compared to supervised and unsupervised methods. To our knowledge, this study made the first effort to jointly exploit the physics-aware deep network and the IFM-inspired semi-supervised learning technique for the UIE task. The proposed method performed better than, or at least comparably to, sixteen baselines on six underwater testing sets in different evaluations. The promising results should be due to the fact that our method is able to not only estimate degradation parameters but also learn the characteristics of diverse underwater scenes.